OpenStudy (anonymous):

How do you factor with negative exponents? I have a couple of problems to do and none of them make any sense to me

Nnesha (nnesha):

like ?

OpenStudy (anonymous):

\[x^2(x-1)^{-2}-4(x-1)^{-1}\]

Nnesha (nnesha):

when you see negative exponent you should flip the fraction \[\huge\rm \frac{ x^{-m} }{1 }=\frac{ 1 }{ x^m }\]

OpenStudy (anonymous):

I understand but how could you factor it if it's a fraction?

OpenStudy (anonymous):

?

Nnesha (nnesha):

oky sorry was answering the phone \[\huge\rm \frac{ x^2 }{ (x-1)^2 }-\frac{ 4 }{ (x-1) }\]

OpenStudy (anonymous):

No problem

Nnesha (nnesha):

alright now find the common denominator :=)

Nnesha (nnesha):

brb